1- If 75 individuals in a population of 735 have the homozygous dominant genotype for a phenotypic trait, what is the allelic frequency for that trait among those individuals?

a. 0.01

b. 0.46

c. 0.75

d. 0.10

e. 0.32

2- A population of geraniums is made up of 428 individuals and you find that 48 are GG, 375 are Gg, and 5 are gg. You also find that the recessive allelic frequency is 0.38. What is the expected number of homozygous recessive individuals if the population is in Hardy-Weinberg equilibrium?

a. 61

b. 375

c. 48

d. 53

e. 5

3- If q2=0.2 and 2pq=0.5 then p=?

a. 0.25

b. 0.55

c. 0.75

d. 1.00

e. 0.30

4- If p=0.35, q=0.65, and the population is in Hardy-Weinberg equilibrium, what proportion of the population will be heterozygous?

a. 0.33

b. 0.24

c. 0.09

d. 0.46

e. 0.17

5- If p=0.25, then q=?

a. 0.50

b. 0.75

c. 1.00

d. 0.00

e. 0.25

6- In a population of 1200 individuals there are 200 with genotype AA, 600 with genotype Aa and 400 with genotype aa. What is the frequency of a?

a. 0.73

b. 0.58

c. 0.21

d. 0.00

e. 1.00

7- A population of geraniums is made up of 428 individuals and you find that 48 are GG, 375 are Gg, and 5 are gg. You also find that the recessive allelic frequency is 0.38. What is the expected homozygous recessive genotypic frequency if the population is in Hardy-Weinberg equilibrium?

a. 0.11

b. 0.01

c. 0.12

d. 0.14

e. 0.38

8- If p=0.4 and q=0.6 and the population is in Hardy-Weinberg equilibrium, what proportion of the population will be homozygous? (Hint: An individual is homozygous if they are not heterozygous, so figure out the proportion of the population that is heterozygous first.)

a. 0.63

b. 0.16

c. 0.24

d. 0.36

e. 0.52

9- A population of geraniums is made up of 428 individuals and you find that 48 are GG, 375 are Gg, and 5 are gg. You also find that the recessive allelic frequency is 0.38. What is the expected frequency of heterozygotes if the population is in Hardy-Weinberg equilibrium?

a. 0.90

b. 0.11

c. 0.62

d. 0.47

e. 0.01

10- A population of doves is made up of 2596 individuals and you find that 96 are DD, 500 are Dd, and 2000 are dd. What is the genotypic frequency of heterozygotes?

a. 0.23

b. 0.77

c. 0.96

d. 0.19

e. 0.03
Answer #1

1.

AA = 75/735 = 0.1021

A = √0.1021 = 0.319 = 0.32

Option E

2.

g = 0.38

gg = 0.38 × 0.38 = 0.1444

Expected number of progenies with gg = 428 × 0.1444 = 61.8 = 61 approx.

Option A

3.

p^2 + 2pq + q^2 = 1

p^2 + 0.5 + 0.2 = 1

p^2 = 1 - 0.7 = 0.3

Option E

4.

Heterozygotes = 2pq = 2 × 0.35 × 0.65 = 0.455

Option D

Please rate high.